A 3.00 -muF and a 4.00 - muF capacitor are connect Can someone help with this?
Answer #1

According to figure

3 micro farad and 4 micro farad

So,

resultant capacitance of these two = C2

1/C2 = 1/(3*10^-6) + 1/(4*10^-6)

C2 = 1.714285714 micro F

And C2 is parallel with 2 micro farad

So equivalent capacitance = C2 + 2 = 1.714285714 + 2 = 3.714285714 micro farad